Consider the following TR and TC function for a firm operating in a perfect competition

market:

TR = 21Q

TC = 1/3Q3 – 5Q2 + 42Q + 81

A. Determine the profit maximizing or loss minimizing equilibrium level of output?

B. Compute the level of profit or loss at the above equilibrium quantity and comment on

the decision of the firm?


1
Expert's answer
2021-12-20T17:57:45-0500

A.

"TR=TC,"

"\frac 13Q^3-5Q^2+21Q+81=0,"


B.

"(\frac 13Q^3-5Q^2+21Q+81)'=0,"

"Q^2-10Q+21=0,"

"Q=3,"

"TR=63,"

"TC=171."